# Underwater Image Enhancement via Dehazing and Color Restoration

### v1

@@ 4.5 Other Comparisons
To further demonstrate that our proposed method has better color correction and broader applications than other methods, qualitative analyses were conducted on Test-S300 and Test-U90. Each image underwent histogram display, Canny edge detection, and saliency object detection. The saliency object detection method was referenced from [51], and Fig. 8 and Fig. 9 show the corresponding results. For Test-S300 and Test-U90, with the exception of WaterFormer, all methods failed to adequately reproduce the color distribution, gradient distribution, and salient feature distribution of the true/reference images. For example, the UDAformer generated color artifacts on Test-S300, which incurred inaccurate edge detection results. Meanwhile, when applied to Test-U90, MLLE uniformly processed all channels, which introduced gradient noise and diminished the salient features of the image. In summary, WaterFormer exhibited outstanding performance in the UIE task and proves to be the most beneficial in related applications.
@@ 4.6 Ablation Study
The proposed WaterFormer incorporates three major components: the CRB, CFB, and loss of combining ${{L}_{chroma}}$ and ${{L}_{Sobel}}$. To demonstrate the effectiveness of each component, an ablation study was performed on the baseline framework with various combinations of these components. Tab. 3 shows the statistical results on the Test-S300 dataset, where the full model ($V_{5}$) outperformed other variants in terms of SSIM and PSNR scores.
Variants CRB CFB ${{L}_{chroma}}$ ${{L}_{Sobel}}$ SSIM↑ PSNR↑
Base w/o w/o w/o w/o 0.872 26.64
${{V}_{1}}$ w/o w/o w/o 0.876 27.09
${{V}_{2}}$ w/o w/o 0.877 27.26
${{V}_{3}}$ w/o 0.891 28.47
${{V}_{4}}$ w/o 0.901 28.56
${{V}_{5}}$ 0.907 29.40
Table 3: Quantitative analysis of the WaterFormer ablation experiments. The red font indicates the best performance.
Fig. 10 displays the visual results of various variants of WaterFormer. The Base model has basic UIE capabilities but retained significant shortcomings in addressing the color degradation and edge artifacts. The $V_{1}$ model, which incorporates the CRB module, could more delicately restore colors. However, due to insufficient feature fusion, the results of the $V_{1}$ model remained lacking. Therefore, by introducing the CFB module, the predictions of the $V_{2}$ model more closely aligned with the true color distribution. The addition of ${{L}_{chroma}}$ in the $V_{3}$ model further enhanced the consistency between inference results and chromaticity of the real images. However, $V_{1}$, $V_{2}$, and $V_{3}$ continued to exhibit noticeable edge artifacts, and this problem was addressed by the inclusion of ${{L}_{Sobel}}$, as demonstrated in the results of the $V_{4}$ inference. Finally, by combining all components, our WaterFormer ($V_{5}$) yielded the closest results to the real images.
@@
In addition, we will demonstrate the effectiveness of other improvements on WaterFormer in Tab. 4, including replacing the ReLU with the FReLU [39] in the MLP structure and replacing the global residual reconstruction layer with the soft reconstruction layer that was proposed in [13] and optimized according to Eq. 3. References [13, 52] also proposed the SK Fusion structure to fuse features from different stages. Unfortunately, we found that the SK Fusion structure was not suitable for feature fusion in underwater images, possibly because SK Fusion does not effectively utilize the global information among different feature channels in underwater images. We present the ablation experiments on SK Fusion in Tab. 4.
@@
WaterFormer 0.907 29.40
FReLU [39]→ReLU 0.905 29.40
UW Soft Recon→Recon 0.904 29.00
UW Soft Recon→Soft Recon [13] 0.904 28.93
CFB→SK fuison [13, 52] 0.903 28.88
Table 4: Quantitative analysis of the ablation experiments on other improvements in WaterFormer. The red font indicates the best performance.
@@ 5 Conclusion
In this work, we introduced WaterFormer, a versatile method to enhance the quality of underwater images. To tackle challenges such as haziness and color shifts, WaterFormer leverages the Vision Transformer architecture and comprises three key components: the DehazeFormer Block, Color Restoration Block (CRB), and Channel Fusion Block (CFB). The DehazeFormer Block transforms hazy features and extract deep-level features. The CRB enhances colors through channel-level transpose matrices. The CFB facilitates feature fusion based on the importance of global information across different feature channels. We also incorporated the Chromatic Consistency Loss and Sobel Color Loss to ensure color fidelity and preserve fine details. The quantitative evaluations, qualitative evaluations, and ablation experiments show that our proposed network performs significantly better than various methods.
Nevertheless, this study has several limitations. First, our work requires a substantial amount of paired data for training, which poses a significant challenge for practical applications. Second, we did not consider the effects of noise such as environmental noise and camera noise during the image processing. Large-scale datasets that cover and fully consider complex types of image noise will be essential to train a more robust and generalizable model.
In our subsequent work, we will consider incorporating a denoising module into the network to reduce the impact of noise and enhance the robustness of the image correction results. In addition, we will use a contrastive learning strategy to reduce the network dependency on data and improve the generalization performance. The denoising process will be integrated into the end-to-end network as a module instead of a separate preprocessing step. We will also explore the processing or analysis based on enhanced results, such as the underwater image classification and measures of image representability. We believe that the analysis of restored underwater images can provide incremental assistance in future underwater visual-imaging tasks.

### v2

@@ 4.4 Other Comparisons
To further demonstrate that our proposed method has better color correction and broader applications than other methods, qualitative analyses were conducted on Test-S300 and Test-U90. Each image underwent histogram display, Canny edge detection, and saliency object detection. The saliency object detection method was referenced from [43], and Fig. 8 and Fig. 9 show the corresponding results. For Test-S300 and Test-U90, MLLE, HFM, and UDAformer failed to accurately reproduce the color, gradient, and saliency features of the reference images. For example, UDAformer introduced color artifacts in Test-S300, leading to inaccurate edge detection results. MLLE applied uniform processing across all channels in Test-U90, introducing gradient noise and reducing saliency features. Conversely, DM-water, GuidedHybSensUIR, and WaterFormer restored the information more accurately. Upon closer inspection, WaterFormer provided the most precise results. For instance, WaterFormer showed a more accurate color distribution in Test-S300, and it yielded results similar to DM-water in Test-U90, while being closer to the reference image in terms of color histogram and saliency feature detection results. In summary, WaterFormer exhibited outstanding performance in the UIE task and proves to be the most beneficial in related applications.
@@ 4.5 Ablation Study
The proposed WaterFormer incorporates four main components: the network backbone, CRB, CFB, and the loss function formed by ${{L}_{\ell_{1}}}$, ${{L}_{chroma}}$ and ${{L}_{Sobel}}$. To validate the effectiveness of each component, a series of ablation studies were conducted on the baseline framework with different combinations of these components. The Test-S300 was used for the ablation experiments, with each image resized to $256\times 256$. The experiments were divided into four parts: first, WaterFormer was broken down into various variants, each composed of different components, and tested on Test-S300; second, the superiority of the CRB design was further explored, and different attention modules were tested as replacements for CRB; third, the CFB fusion mechanism was visualized, with pixel-wise addition used as a comparison; fourth, other optimizations and experimental parameters of WaterFormer were tested by ablation.
@@
Variants CRB CFB ${{L}_{chroma}}$ ${{L}_{Sobel}}$ SSIM↑ PSNR↑
Base w/o w/o w/o w/o 0.876 27.32
${{V}_{1}}$ w/o w/o w/o 0.890 28.50
${{V}_{2}}$ w/o w/o 0.892 28.87
${{V}_{3}}$ w/o 0.907 30.35
${{V}_{4}}$ w/o 0.913 30.46
${{V}_{5}}$ 0.917 31.18
Table 4: Quantitative analysis of the WaterFormer ablation experiments. The bold font indicates the best performance.
Variants Evaluation. Fig. 10 and Fig. 4 display the statistical results and visualization results on the Test-S300. The base model was composed only of the backbone and was trained using the ${{L}_{\ell_{1}}}$. It had basic UIE capabilities but still retained significant shortcomings in addressing color degradation and edge artifacts. The $V_{1}$ model introduced the CRB, which enabled it to recover color more accurately. However, due to insufficient integration of different features, the results of $V_{1}$ still exhibited a large amount of color artifacts. Therefore, by introducing the CFB, the predictions of the $V_{2}$ model exhibited fewer color artifacts and a more realistic color distribution. The addition of ${{L}_{chroma}}$ to the $V_{3}$ model further improved the consistency between inference results and the chromaticity of real images. However, $V_{1}$, $V_{2}$, and $V_{3}$ continued to exhibit noticeable edge artifacts, such as the white rim of the bucket, and this problem was addressed by including ${{L}_{Sobel}}$, as demonstrated in the results of the inference $V_{4}$. Finally, by combining all components, our WaterFormer ($V_{5}$) yielded the closest results to the real images with the highest SSIM and PSNR scores.
CRB Evaluation. The design of CRB is based on the mechanism of self-attention between channels, as proposed by  [34] and  [35]. CRB extends this mechanism and applies it effectively to the UIE task. To visualize the effect of CRB,  11 shows the output of the feature map of the first 2-block CRB in the $V_{1}$ variant. For comparison, the feature map output of the first 2-block DehazeFormer Block in the Base model is included. We used maximum projection and the jet color map to plot the heatmap, where blue represents lower attention from the network, and red represents higher attention. In Test-S300, we selected two representative types of water for validation. The first row represents clear water, with only color degradation, but clear foreground and background. The second row represents turbid water, with both color degradation and severe occlusion. As seen in Fig. 11, the DehazeFormer Block focuses on the foreground, and its performance decreases significantly when most of the foreground is occluded. In contrast, CRB can focus on global color correction, and its heatmap more accurately reflects the distribution of color bias. This shows that the DehazeFormer Block and CRB handle different subtasks in the UIE task, with CRB assisting the DehazeFormer Block in feature extraction while performing color restoration. Furthermore, to demonstrate the higher performance brought about by CRB, we replaced MTDA in  [34] and MCC in [35] with CRB. The results are shown in Tab. 5. Additionally, RLN in the CRB significantly improved color restoration, and for comparison, BatchNorm and LayerNorm were included in the ablation study. To further highlight the benefits of CRB and channel self-attention, we also included popular attention mechanisms, SE Block [44] and CBAM [45], in the ablation experiment.
[…]
Fig. 11: Visualization of the CRB working mechanism. (a) Input, (b) result of DehazeFormer Block, (c) result of CRB, (d) GT.
@@
WaterFormer 0.917 31.18
CRB→MTDA [34] 0.910 30.01
CRB→MCC [35] 0.912 30.43
CRB→SE Block [44] 0.909 30.00
CRB→CBAM [45] 0.907 29.89
RLN→BatchNorm 0.916 31.05
RLN→LayerNorm 0.915 30.85
Table 5: Quantitative analysis of CRB ablation experiments. The bold font indicates the best performance.
CFB Evaluation. To better validate the effectiveness of the CFB, its fusion mechanism was visualized and compared with pixel-wise addition, as shown in Fig. 12. The first CFB example is displayed, where its two input feature maps come from the image embedding layer (composed of a convolutional layer with a 3x3 kernel size) and the 2-block CRB. We also selected clear and turbid water types for the experiment in Test-S300, which are presented in the first and second rows, respectively. In Fig. 12, the image embedding layer can only extract shallow features, and its heatmap reflects mainly the pixel distribution of the input. On the other hand, as described in the previous subsection, the heatmap generated by CRB displays the distribution of color bias information. From the heatmap, it can be seen that if the two feature maps are added pixel-wise, many details are lost, and the WaterFormer inference results are therefore affected. However, after fusion using the CFB, the results for both clear and turbid water are presented with more distinct layers and better detail preservation (as indicated by the purple arrow).
[…]
Fig. 12: Visualization of the CFB fusion mechanism. (a) Input, (b) result of image embedding layer, (c) result of CRB, (d) result of CFB, (e) result of addition, (f) GT.
Other Evaluation. The effectiveness of other improvements in WaterFormer was demonstrated in Tab. 6, including the replacement of the ReLU with the FReLU [33] in the MLP structure and the replacement of the global residual reconstruction layer with the soft reconstruction layer proposed in [14] and optimized according to Eq. 3. The study on the size of the sliding window in ${{L}_{chroma}}$ can be found in Tab. 7. The study on the scaling coefficients of each loss term in the loss function can be found in Tab. 8.
Variants SSIM↑ PSNR↑
WaterFormer 0.917 31.18
FReLU [33]→ReLU 0.915 30.82
UW Soft Recon→Recon 0.911 30.26
UW Soft Recon→Soft Recon [14] 0.916 30.79
Table 6: Quantitative analysis of the ablation experiments on other improvements in WaterFormer. The bold font indicates the best performance.
Sliding window size SSIM↑ PSNR↑
$11\times 1$ 11 0.916 31.12
$13\times 13$ 0.916 31.04
$15\times 15$ 0.917 31.18
Table 7: Quantitative analysis of the ablation experiments on size of the sliding window in ${{L}_{chroma}}$. The bold font indicates the best performance.
${\lambda}_{1}$ ${\lambda}_{2}$ ${\lambda}_{3}$ SSIM↑ PSNR↑
1 1 1 0.912 30.53
2 1 1 0.909 30.03
1 2 1 0.915 30.88
1 1 2 0.917 31.18
2 2 1 0.912 30.48
2 1 2 0.912 30.84
1 2 2 0.916 31.29
2 2 2 0.911 30.20
Table 8: Quantitative analysis of the ablation experiments on scaling coefficients of each loss term in the loss function. The bold font indicates the best performance.
@@ 5 Conclusion
In this work, we introduce WaterFormer, a versatile method to enhance the quality of underwater images. To address challenges such as haziness and color shifts, WaterFormer leverages the ViT architecture and comprises three key components: the DehazeFormer Block, the Color Restoration Block (CRB) and the Channel Fusion Block (CFB). The DehazeFormer Block transforms hazy features and extracts deep-level features. The CRB enhances colors through channel-level transpose matrices. The CFB facilitates feature fusion based on the importance of global information across different feature channels. We also incorporated the Chromatic Consistency Loss and Sobel Color Loss to ensure color fidelity and preserve fine details. Quantitative evaluations, qualitative evaluations, and ablation experiments show that our proposed network performs significantly better than various methods.
However, this study has several limitations. First, our work requires a substantial amount of paired data for training, which poses a significant challenge for practical applications. Second, we did not consider the effects of noise, such as environmental noise and camera noise, during image processing. Large-scale datasets that cover and fully consider complex types of image noise will be essential to train a more robust and generalizable model.
In our subsequent work, we will consider incorporating a denoising module into the network to reduce the impact of noise and enhance the robustness of the image correction results. In addition, we will use a contrastive learning strategy to reduce the dependency of the network on data and improve generalization performance. The denoising process will be integrated into the end-to-end network as a module instead of a separate preprocessing step. We will also explore the processing or analysis based on enhanced results, such as the underwater image classification and measures of image representability. We believe that the analysis of restored underwater images can provide incremental assistance in future underwater visual-imaging tasks.